OpenStudy (anonymous):

without drawing the graph of the equation, determine how many points the parabola has in common with the x-axis and whether its vertex lies above, on, or below the x-axis. y=4x^2-12x+12 A. 2 points in common; vertex below x-axis B. no points in common; vertex above x-axis C. 1 point in common; vertex on x-axis D. 2 points in common; vertex above x-axis

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

Calculate the discriminant.

OpenStudy (anonymous):

x^2 - x + 2 = 0 the discriminant is b²-4ac when you have ax²+bx+c therefore, you have a=1, b=-1, c=2, and b²-4ac = (-1)² - 4(1)(2) = 1 - 8 = -7

OpenStudy (anonymous):

make since?

OpenStudy (anonymous):

thats a exsample just to help you understand. here let me solve for your problem

OpenStudy (anonymous):

Without drawing the graph of the equation, determine how many points the parabola has in common with the x-axis and whether its vertex lies above, on, or below the x-axis. y = 4x^2 - 12x + 12 y = 4(x² - 3x) + 12 y = 4(x² - 3x + 3/4) + 12 - 3 y = 4(x - 3/2)² + 9 vertex: (3/2, 9) no points in common; vertex above x-axis

OpenStudy (anonymous):

So your answer would be B

OpenStudy (anonymous):

does this make since @yayah2001 ?

OpenStudy (anonymous):

thank you, I understand a lot better now.

OpenStudy (anonymous):

Np! :D